OpenStudy (anonymous):

One of the main components of hard water is CaCO3. When hard water evaporates, some of the CaCO3 is left behind as a white mineral deposit. If a hard water solution is saturated with CaCO3, what volume of the solution has to evaporate ot deposit 1*10^{2} mg of CaCO3? I do not know what to do here, but I know this is a solubility equilibria problem.

OpenStudy (jfraser):

This is a solubility equilibrium, so to solve this you need the KSP of CaCO3 to know what the concentration of Ca and CO3 ions is in order to evaporate the proper volume of solution

OpenStudy (anonymous):

And then I just use dimensional analysis and find the volume?

OpenStudy (anonymous):

Mos to the rescue! Yes.

OpenStudy (mos1635):

set C consetration of saturated solution CaCO3->Ca(2+) + CO3(2-) C - - -C +C +C - C C IN STURATED CONDITION Ksp=[Ca2+]*[CO3 2-] Ksp=C^2 C=sqr Ksp.. now C=n/V n=C*V m/Mr=C*V V=m/(Mr*C) note convert mgr to gr

OpenStudy (mos1635):

is Ksp given??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Hey Mos, I actually got the answer for this question. The question I really needed help with was the following: Cu(s)|CuI(s)|I-(aq, 1.0M) || Cu2+(aq, 1.0M) | Cu(s) Find Ecell for this reaciton, and the Ksp is given.

OpenStudy (anonymous):

For the above line notation, I understand that the reduction r'xn is this: Cu2+(aq) + 2e- --> Cu(s) But the oxidation reaction does not make sense to me, from what it is showing.

OpenStudy (anonymous):

Apparently, it shows that: CuI(s) --> Cu(s) + I-(aq) + 1e- <-- How can that be? I was thinking: CuI(s) ---> Cu2+(aq) + I-(aq) + 1e-

OpenStudy (anonymous):

What is the Cu(s) on the LHS of the notation there for?

OpenStudy (mos1635):

sorry not familiar with this way of given data.....

OpenStudy (anonymous):

OH! Is the Cu(s) just indicating to you what the electrode is? But then, it would still not make sense why they did not include Cu2+(aq) on the LHS of the line notation..